List of participants

Sergey Alekhin  IHEP Protvino  alekhin@sirius.ihep.su
Elke-Caroline Aschenauer DESY Hamburg  elke.aschenauer@desy.de
Daniel Boer  U Amsterdam  dboer@nat.vu.nl

Yuraj Bracinik  MPI Munich  juraj.bracinik@desy.de

Gerd Buschhorn  MPI Munich  gwb@mppmu.mpg.de
Allen Caldwell  MPI Munich  caldwell@mppmu.mpg.de
Vladimir Chekelian  MPI Munich  shekeln@mail.desy.de
Francois Corriveau  U McGill  francois@mail.desy.de
Jaroslav Cvach  Academy of Science Prague cvach@fzu.cz
Kai Diener  PSI Villigen  kai.diener@psi.ch
Elisabetta Gallo  U Florence, INFN  gallo@mail.desy.de
Claudia Glasman  U Madrid  claudia@mail.desy.de
Günter Grindhammer  MPI Munich  guenterg@desy.de
Gösta Gustafson  U Lund  gosta@thep.lu.se

Gudrun Heinrich  Durham, IPPP  gudrun.heinrich@durham.ac.uk
Leif Jönsson  U Lund  leif.jonsson@hep.lu.se

Rosita Jurgeleit  MPI Munich  roj@mppmu.mpg.de
Dorian Kcira  U Wisconsin  kcira@mail.desy.de
Christian Kiesling  MPI Munich  cmk@mppmu.mpg.de
Max Klein  DESY Zeuthen  klein@ifh.de

Bernd Kniehl  U Hamburg  kniehl@mail.desy.de

Gustav Kramer  U Hamburg  kramer@mail.desy.de
Sergey Levonian  DESY Hamburg  levonian@mail.desy.de
Markos Maniatis  U Hamburg  markos@mail.desy.de
Sven Moch  DESY Zeuthen  moch@ifh.de
Leszek Motyka  Jagellonian U Cracow  leszekm@th.if.uj.edu.pl

Otto Nachtmann  U Heidelberg  o.nachtmann@thphys.uni-heidelberg.de

Wolfgang Ochs  MPI Munich  wwo@mppmu.mpg.de
Sanjay Padhi  DESY Hamburg  spadhi@mail.desy.de
Robert Peschanski  CNRS Saclay  pesch@spht.saclay.cea.fr
Ringaile Placakyte  MPI Munich  ringaile@mppmu.mpg.de
Benjamin Portheault  LAL Orsay  portheau@lal.in2p3.fr
Jon Pumplin  Michigan State U  pumplin@pa.msu.edu
Zuzana Rurikova  MPI Hamburg  zuzana.rurikova@desy.de
Ingo Schienbein  DESY Hamburg  schien@mail.desy.de
Fridger Schrempp  DESY Hamburg  fridger.schrempp@desy.de
Uta Stösslein  DESY Hamburg  uta.stoesslein@desy.de
Paul Thompson  U Birmingham  thompspd@mail.desy.de

Michael Tytgat  U Gent  michael@inwfsun1.UGent.be

Nikolai Vlasov  U Freiburg  vlasov@mail.desy.de
Biljana Vujicic  MPI Munich  bivuj@mppmu.mpg.de

Rik Yoshida  Argonne  rik.yoshida@desy.de